A lever is used to lift a load of 320Kg. The effort applied is 210 N and it moves through a distance of 1.7 m. The load moves 345mm


Determine the following:


  1. Mechanical Advantage
  2. Displacement ratio
1
Expert's answer
2021-03-10T17:14:16-0500

1)

"MA=\frac{F_2}{F_1}=\frac{mg}{F_1}=\frac{320\cdot 9.8}{210}=15,"

2)

"DR=\frac{d_1}{d_2}=\frac{1.7}{0.345}=4.9."